Find the lateral area of the small cone that has a radius of 3 inches and a height of 4 inches.
Answer by pushpaharan(47)

Surface area of cone is = (PI)*R*S + (PI)*R^2, where R- radius, S- Slant height.
Slant height can be found using Pythagorean Theorem, where you know the height, and radius. Then slant height will be hypotenuse.
there fore s= sqrt( h^2 + r^2) = 5
Surface area = 75.43
If it doesn't haVE CIRCULER LID, THEN NO NEED TO CONSIDER (PI)*R^2